DK4L:

Three polynomials are factored below, but some coefficients and constants are missing. All of the missing values of a, b, c, and d are integers. 1. x^2-8x+15=(ax+b)(cx+d) 2. 2x^3-8x^2-24x=2x(ax+b)(cx+d) 3. 6x^2+14x+4=(ax+b)(cx+d) Fill in the table with the missing values of a, b, c, and d.

DK4L:

How do u get the answer?

jhonyy9:

1. x^2 -8x +15 = you need get two numbers with sum equal -8 and with product equal +15

DK4L:

ok

jhonyy9:

what are these numbers ?

DK4L:

I got the question right

jhonyy9:

ok what are these two numbers ?

DK4L:

idk i guess

jhonyy9:

ok for the first one -8 = -5 - 3 bc. (-5)*(-3) = +15

DK4L:

Ok but for the ist one I got (1, -5, 1, 3)

jhonyy9:

this is your first equation 1. x^2-8x+15= not what is posted in the first answer

jhonyy9:

1. x^2-8x+15=(ax+b)(cx+d) a = 1 b = -3 c = -5

DK4L:

oops i posted the wrong equation its x^2+2x-8=(ax+b)(cx+d)

jhonyy9:

indifferent the way to solve it is the same

DK4L:

so this one would be a=1 b=-2 c=1 d=4

jhonyy9:

correct

DK4L:

ok thnx

jhonyy9:

np yw

DK4L:

hey since your on level 99 do u move up to level 100